OpenStudy (anonymous):

Confirm that f and g are inverses by showing that f(g(x)) = x and g(f(x)) = x. f(x)= 4/x and g(x)= 4/x If someone could please explain that'd be great :)

OpenStudy (anonymous):

f (g(x) ), let x=g(x) f(4/x) = (4/x) f(4/x) = 4/(4/x) = x

OpenStudy (anonymous):

@larryboxaplenty So then the next part would be g(f(x)), let x=f(x) g(4/x) = (4/x) g(4/x)= 4/(4/x)=x I understand the math I just don't know exactly what the question is asking and why am I solving the equation like this.